A starting player for a major college basketball team made only 5% of her free throws last season. During the off season she worked on developing a softer shot in the hope of improving her free throw accuracy. In the first eight games this season she made 22 out of the 42 free throws she attempted. Did she improve her accuracy?
Answer #1

Ans:

H0:p=0.5

Ha:p>0.5

Sample proportion=22/42=0.5238

Test statistic:

z=(0.5238-0.5)/sqrt(0.5*(1-0.5)/42)

z=0.31

p-value=P(z>0.31)=0.378

Fail to reject the null hypothesis.

There is not sufficient evidence to conclude that she improve the accuracy.